OpenStudy (anonymous):

Find the general equation of the line Containing the points (-9,5) and (3, 6)

OpenStudy (anonymous):

um i think i know this (google search engine)

OpenStudy (anonymous):

i cant figure it out.

OpenStudy (anonymous):

ok ill help

OpenStudy (anonymous):

it has to be in Ax + By = C form and the answer i put was wrong

OpenStudy (luigi0210):

Find the slope first. \(\Large m=\frac{y_{2}-y_{1}}{x_{2}-x{1}}\)

OpenStudy (anonymous):

i did that. i got -1/2

OpenStudy (anonymous):

my teacher is the worst. he doesnt know how to do math lol and just confuses every one.

OpenStudy (luigi0210):

\(\Large m=\frac{6-5}{3-(-9)}\) So should be \(\Large m=\frac{1}{12}\)

OpenStudy (anonymous):

ohhhhh that explains alot

OpenStudy (anonymous):

thank u!

OpenStudy (luigi0210):

Yup :) You can finish it from here right?

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

so in the y=mx+b i would choose any equation right

OpenStudy (anonymous):

for example the (3,6) to replace the x and y values in that eq'n

OpenStudy (anonymous):

so it would be 6=(1/12)(3)+b 6=0.25+b 5.75=b right?

OpenStudy (luigi0210):

That's a pretty awesome method o: but yes, that would be correct :) So your equation would be: \(\Large \color{blue}{y=\frac{1}{12}x+5.75}\) ^That would be it in slope intercept form

OpenStudy (anonymous):

ok and to get it into Ax+By=C it would be (1/12)x-y=-5.75 since x has to be positive right

OpenStudy (anonymous):

omg im gonna flip out .. im sending a screenshot

OpenStudy (luigi0210):

Yup, but when doing Standard form, there can't be any fractions or decimals in front of the A. So you have to multiply everything by 12 to get rid of that 1/12 \(\Large \color{red}{12}(\color{blue}{\frac{1}{12}x-y=-5.75)}\) So it would just be \(\Large \color{green}{x-12y=-69}\) Well ain't that a pretty number <.<

OpenStudy (luigi0210):

And is everything alright? o.O

OpenStudy (anonymous):

sorry that took awhile

OpenStudy (anonymous):

true thank u .. this is a tough subject for me.. i thought stats would be easy :( ty so so much

OpenStudy (luigi0210):

It's alright, I understand :P And if you need anymore help just PM me or post another question, since the site isn't working properly that'd be best ~ And good luck \(\Large \color{red}{\ddot\smile}\)